4 FeS2+15 O2>2Fe2 O3+8SO3
How many moles of O2 would be required to completely react with 2.75 moles of FeS2?
How many moles of Fe2O3 would be theoretically produced by the conditions in the previous question above? How many moles of SO3?
Can someone please show me the work along with the answer?
1
Expert's answer
2020-06-15T14:43:58-0400

Explanation:

Write the balanced equation as presented.

4FeS2+15O2→2Fe2O3+8SO3

i)moles of O2

4mole FeS2=2.75moles FeS2

15 Mole O2=2.75*15/4=10.31 moles O2

ii)Moles of Fe2O3?

4 Mole FeS2=2Mole FeO3

But, 4 Mole FeS2=2,75 mole FeS2

2 mole FeO3=2*2.75/4=1.375 Moles Fe2O3

iii)Moles of So3?

4 mol FeS2=2.75 Mol FeS2

8mol So3=8*2.75/4=5.5 mol So3